OpenStudy (anonymous):

Tub A contains 23% apple cider. Tub B contains 65% of apple cider. How many ounces of cider from each tub must be gotten to make a 20 oz. drink that is 35% apple cider

OpenStudy (mertsj):

x= ounces of 23% 20-x= ounces of 65% .23x+.65(20-x)=.35(20)

OpenStudy (mertsj):

Multiply by 100 to remove the decimal points and solve.

OpenStudy (mertsj):

\[23x+1300-65x=700\]

OpenStudy (mertsj):

-42x=-600 x=14.3 ounces of 23% 20-x= 5.7 ounces of 65%